OpenStudy (anonymous):

Which expression is equivalent to (–2)(a + 6)

OpenStudy (mathmale):

apparently there are several possible answers to choose from. Would you please post them?

OpenStudy (solomonzelman):

Lets say you had the following expressions: 3(b-4) the way to expand it is.... 3(b-4) = 3 × b + 3 × (-4) = 3b +(-12) = 3b -12. Can you do (–2)(a + 6) the same way ?

OpenStudy (anonymous):

sure ehh maybe

OpenStudy (solomonzelman):

Try to do it... type what you do in here, so that I can correct you, if you get any of the process wrong.

OpenStudy (solomonzelman):

I am going to start you off... (–2)(a + 6) = (–2) × a + (-2) × 6 = (and you please continue....)